OpenStudy (amonoconnor):

Hey @ganeshie8! I have another (hopefully quicker) question. Can you check this: ORIGINAL: "(x^3) - (1/(x^3))" MY ANT.DRVTV. = ((x^4)/4) + (1/(2x^2)) + C (?) :)

ganeshie8 (ganeshie8):

Correct ! good job !!

OpenStudy (amonoconnor):

YAYAYAY!

ganeshie8 (ganeshie8):

Also for verifying the answer, u may use the wolfram : http://www.wolframalpha.com/input/?i=%5Cint+%28x%5E3%29+-+%281%2F%28x%5E3%29%29 its a good site for checking answers :)

OpenStudy (amonoconnor):

Lols, I get the hint ;) Alrighty, I'll check it out.

OpenStudy (whpalmer4):

Taking the derivative of your answer and comparing it with what you started with is not bad, either, and gives you extra practice doing derivatives :-)

ganeshie8 (ganeshie8):

thats a very good idea actually, instead of dumbly checking in wolfram :)

ganeshie8 (ganeshie8):

take the derivative of ur answer, u should get back the function u started wid

OpenStudy (whpalmer4):

there are cases where you may end up with something that is equivalent but expressed differently, say in trig stuff, but then go ahead and use wolfram alpha :-)

OpenStudy (amonoconnor):

Ahhhhh... That's awesome! I didn't put two and two together until you mentioned that... That's brilliance. :))

OpenStudy (amonoconnor):

Haha, okay. Good deal :)

OpenStudy (isaiah.feynman):

Wolfram is the lazy way. :P

OpenStudy (whpalmer4):

but it's like multiplying out the result to check your factoring work: it helps you recognize the patterns so the next time you might not even have to do the work, or as much of it.

OpenStudy (amonoconnor):

Lols. Hating a bit are we ;)

OpenStudy (amonoconnor):

Thank you to all who positively contributed! You were immensely helpful. :)

OpenStudy (whpalmer4):

I'm not hating, I run Mathematica (desktop version of Wolfram Alpha, you might say) all the time. But I try to do enough of the work by hand that I keep the skills sharp, and use the big M to keep me honest or the real drudge work. I'm all for having computers do the drudge work, as long as you're still able to do it yourself!

OpenStudy (amonoconnor):

Lols, sorry. I was talking to Isaiah, haha. :) sorry, sorry, sorry!!

OpenStudy (amonoconnor):

I didn't hit enter before you posted under him.

ganeshie8 (ganeshie8):

hahah happens xD 'lazy' is very much a positive word... u knw there is a theory that says somthing like : all efficient folks are lazy...

ganeshie8 (ganeshie8):

“Laziness is the first step towards efficiency.” ― Patrick Bennett u wil get to hear these kindof stuff more if u ever enter into computer science careers...

ganeshie8 (ganeshie8):

lol jk ;)

OpenStudy (amonoconnor):

Lols, nice.... ;)

OpenStudy (whpalmer4):

or as GB Shaw said, "The reasonable man adapts himself to the world. The unreasonable man persists in trying to adapt the world to himself. All progress, therefore, depends upon the unreasonable man."

ganeshie8 (ganeshie8):

lol nice one xD all inventors/scientists are unreasonables :P

OpenStudy (amonoconnor):

Oh God... Turned the page: mind-stoppage. Okay, so my next one is: "piCOSpiX"

OpenStudy (amonoconnor):

Ick, that looks bad. Umm, Here: "(pi)(cos(pi)x)" ...O believe is how that would be broken up (?)

OpenStudy (amonoconnor):

*I believe...

ganeshie8 (ganeshie8):

heard of 'substitution method ' before ?

ganeshie8 (ganeshie8):

\(\large \int \pi \cos(\pi x) dx\)

OpenStudy (amonoconnor):

Is it related to the Chain Rule?

ganeshie8 (ganeshie8):

yup !!

ganeshie8 (ganeshie8):

or u can use advanced guessing if u want

OpenStudy (amonoconnor):

I haven't seen that before, but okay, I'm familiar with the Chain Rule.

OpenStudy (amonoconnor):

. . . So, I'm assuming I should use form above. What's the first step?

ganeshie8 (ganeshie8):

substitute, \(\pi x = t \) differentiate both sides : \(\pi dx = dt \) solve for \(dx\) \(dx = \frac{dt}{\pi}\)

ganeshie8 (ganeshie8):

so the integral becomes : \(\large \int \pi \cos(t) \frac{ dt}{\pi} \) \(\large \int \cos(t) dt \)

OpenStudy (amonoconnor):

Where did the "t", and the "dt/pi" come from... (?)

ganeshie8 (ganeshie8):

integral of \(cos(t)\) is simply \(\sin(t) + c\) plug back the t value, the final answer wud be : \(\sin(\pi x) + c\)

ganeshie8 (ganeshie8):

thats one way, its called 'substitution method'

OpenStudy (amonoconnor):

So... "t" represents whatever 'value' is in place of the theta value, representing it as a term?

ganeshie8 (ganeshie8):

yes, we substituted \(\pi x = t\)

OpenStudy (amonoconnor):

So... What happens to the first pi, before the cos?

ganeshie8 (ganeshie8):

good question :) pi is just a constant. so it stays same.

OpenStudy (amonoconnor):

So... To find the anti derivative of this, we take the derivative of the whole thing [ =(pi(sin(pix)) ] Times d(theta), or dt ? So... Is it: pi(SIN(pi))

OpenStudy (amonoconnor):

Whoops, "+ C"

OpenStudy (amonoconnor):

? Am I even remotely on the right track? :/

ganeshie8 (ganeshie8):

yes you're on right track, but lets do it from scratch again.. to get u use to 'substitution method' :)

OpenStudy (amonoconnor):

Okay.

ganeshie8 (ganeshie8):

first step : substitute \(\pi x = t\), and find \(dx\) in terms of \(dt\)

ganeshie8 (ganeshie8):

\(\pi x =t \) differentiate both sides, wat do u get ?

OpenStudy (isaiah.feynman):